A pharmacist is to prepare 15 liters of eye drops. eye drop solution must have a 4% active ingredient, but the pharmacist only has 10% solution and 2% solution in stock. How much of the 10% soulution should be used to fill the prescription?
Answer by stanbon(75887)

A pharmacist is to prepare 15 liters of eye drops. eye drop solution must have a 4% active ingredient, but the pharmacist only has 10% solution and 2% solution in stock. How much of the 10% soulution should be used to fill the prescription?
------------------------
Equation:
active + active = active
0.10x + 0.02(15-x) = 0.04*15
Multiply thru by 100 to get:
10x + 2*15 - 2x = 4*15
10x = 2*15
x = 3 liters (amount of 10% solution needed in the mixture)
=========================
Cheers,
Stan H.